Roseland Company purchased merchandise from Royal Corp. on September 30, year 1. Payment was made in the form of a noninterest-bearing note requiring Roseland to make six annual payments of $4,400 on each September 30, beginning on September 30, year 4.

Required:

Calculate the amount at which Roseland should record the note payable and corresponding purchases on September 30, year 1, assuming that an interest rate of 12% properly reflects the time value of money in this situation.
Answer #1

Compute amount at which R should record the note payable as shown below: J 1 2 3 4 34 5 6 36 7 37 8 38 9 39 10 40 Total K L M

Calculation formulas: AJ K L M 29 Year PV factor Payment Note payable 30 1 =1/1.124330 31 2 =1/1.124131 32 3 =1/1.12AJ32 33 4